# ChemKANs for Combustion Chemistry Modeling and Acceleration

### v1

@@ Abstract
Efficient chemical kinetic model inference and application for combustion problems is challenging due to large ODE systems and wideley separated time scales. Machine learning techniques have been proposed to streamline these models, though strong nonlinearity and numerical stiffness combined with noisy data sources makes their application challenging. The recently developed Kolmogorov-Arnold Networks (KANs) and KAN ordinary differential equations (KAN-ODEs) have been demonstrated as powerful tools for scientific applications thanks to their rapid neural scaling, improved interpretability, and smooth activation functions. Here, we develop ChemKANs by augmenting the KAN-ODE framework with physical knowledge of the flow of information through the relevant kinetic and thermodynamic laws, as well as an elemental conservation loss term. This novel framework encodes strong inductive bias that enables streamlined training and higher accuracy predictions, while facilitating parameter sparsity through full sharing of information across all inputs and outputs. In a model inference investigation, we find that ChemKANs exhibit no overfitting or model degradation when tasked with extracting predictive models from data that is both sparse and noisy, a task that a standard DeepONet struggles to accomplish. Next, we find that a remarkably parameter-lean ChemKAN (only 344 parameters) can accurately represent hydrogen combustion chemistry, providing a 2 acceleration over the detailed chemistry in a solver that is generalizable to larger-scale turbulent flow simulations. These demonstrations indicate potential for ChemKANs in combustion physics and chemical kinetics, and demonstrate the scalability of generic KAN-ODEs in significantly larger and more numerically challenging problems than previously studied.
keywords:
Kolmogorov-Arnold Networks, Chemical Kinetics, Dynamical Systems , Reacting Flows , Model Inference , Computational Acceleration1 Introduction
Chemical kinetic modeling is a key backbone of many areas of science and engineering, from biological systems to energy conversion. Traditional hurdles to effective kinetic modeling for combustion have included large and complex reaction mechanisms that can be challenging to derive manually via expert knowledge [1], and the substantial size and stiffness of these chemical kinetic systems [2, 3] leading to large computational cost when integrating their solutions. Recent developments in machine learning have shown promise in helping to address these two issues [4] through streamlined model inference techniques and efficient, data-driven solvers.
@@
Machine learning tools have been proposed using diverse algorithms, presumed model forms, and optimization approaches to accomplish automatic kinetic model discovery from various types of experimental data. The Chemical Reaction Neural Network approach [5], for example, is capable of inferring reaction networks and parameters from limited species trajectory or heat release data [6, 7, 8] by directly enforcing the Arrhenius and mass action laws in a neural network structure. The Sparse Identification of Nonlinear Dynamics (SINDy) approach is similarly capable of extracting models from experimental data by assuming various functional relationship building blocks and learning the precise forms needed to fit the data [9]. Further optimization and inverse modeling tools exist for other chemical kinetic inference problems [10, 11, 12], with a key piece of many physics-based model inference techniques being a certain (and often substantial) degree of prior knowledge of the governing equations, reaction pathways, and reactants. A critical capability required of such kinetic model discovery algorithms, especially in the field of combustion, is robustness to noisy data or uncertain model pathways and parameters [5, 13, 7], with realistic models often containing significant uncertainty on the kinetic parameters.
@@
On the solver side, dimension reduction [6, 14] and computational acceleration [15, 16, 17, 18] are two general tools that aim to address the issues of large mechanisms and slow stiff solvers. For instance, Owoyele and Pal [17] recently proposed ChemNODE, a creative and high-performing tool which leverages the neural ODE concept of Chen et al. [19] to replace a complete chemical kinetic model with a collection of neural networks, one for each tracked thermochemical quantity. By directly linking the current thermochemical state to the chemical source term with no other problem-specific treatment, computational acceleration was enabled in the as-studied homogeneous reactor while retaining the generalizability of the surrogate model to higher-dimensional reacting flows where such acceleration becomes significantly more meaningful. Other recent works leverage DeepONets [20] to directly learn stiff integrators using neural operators, either with problem-specific network structures [21, 14] or by mapping from the current state to the source term, like in ChemNODE [22, 18], allowing for significant computational acceleration downstream. These strengths all come with drawbacks, however. Owoyele and Pal [17], for example, found that while the neural ODE approach’s clever exploitation of the dynamical structure of chemical kinetic models can provide high accuracy, the nonlinearity inherent to such models creates a challenging inference problem for the underlying MLP layers that led the authors to omit a handful of species (including the key H radical) and break the training up into multiple unique and likely redundant networks, rather than a single cohesive architecture. DeepONet techniques are cheap to evaluate and capture steady state behavior well, but their accuracy can suffer in stiff regions of the data that the integrator (which with DeepONets must be inferred directly by the network, as they do not explicitly leverage existing ODE solvers as Neural ODE approaches do) can learn to skip over without significant penalties. We thus find that despite these recent novel and productive efforts, the training of efficient surrogate models for combustion chemistry remains a challenging task with open questions due to stiff behavior in the solution profiles and numerical instability in the nonlinear training processes.
@@
Kolmogorov-Arnold Networks were proposed recently [23] as an alternative to multi-layer perceptrons for general neural network applications, where instead of learning weights and biases on fixed activation functions, the shapes and magnitudes of the activation functions themselves are learned via gridded basis function sums and products. This shift was proposed to increase neural convergence rates, accuracy, and generalization. Echoing the development of traditional MLPs, physics-informed KAN structures were proposed shortly after, where certain knowledge of physical laws can be embedded in the training process to help the KAN converge to a physically meaningful solution [24, 25, 26]. The inference benefits of KANs were additionally demonstrated to extend to dynamical system modeling in the Kolmorogorov-Arnold Network Ordinary Differential Equations (KAN-ODEs) framework [27], where KANs replaced MLPs in the neural ODE algorithm [19], and have since been demonstrated in a variety of settings, including predator-prey dynamics, shock formation, complex equations, phase separation [27], personalized cancer treatment [28], and flashover prediction [29]. In short, a KAN network serves as the gradient getter, while a standard ODE solver integrates the solution profiles as in a traditional numerical solver. KAN-ODEs were shown to retain all major KAN benefits while also accessing the dynamical system inference capabilities of the neural ODE framework, which would appear to lend them to efficient chemical kinetic system modeling. However, a few key questions remain. KAN-ODEs have so far only been tested in relatively small systems (up to two-dimensional state variables), making their applicability and performance in larger, practical combustion systems unknown. Additionally, KANs in general have been shown to suffer substantially when trained with noisy datasets [30]. While we theorize that their direct coupling to ODE integrators combined with their sparse parameterization and smooth activations should provide KAN-ODEs with strong robustness to noise regardless of previously reported issues in generic KANs, this has not yet been studied quantitatively.
@@
In this work, we aim to develop a chemistry KAN-ODE (ChemKAN) framework for chemical kinetic modeling by designing the Kolmorogorov-Arnold network gradient getter to learn the relationship between the current thermochemical state and the chemical source terms. Developed across two case studies of increasing complexity, the ChemKAN framework contains a physics-informed, two-stage training process that enforces the direct coupling between species production and heat release, and additionally contains a soft constraint for species element conservation. We further stabilize the optimization problems by implementing forward sensitivity analysis. Stiff chemistry is fully resolved via an attached numerical ODE solver. We study two cases here to explicitly probe the key behaviors and gaps in the model inference and solver acceleration literature identified above, both of which are addressed with ChemKANs. First, we demonstrate the capability of ChemKANs to extract realistic and multi-species models from synthesized experimental data in a comparison against DeepONets (DONs) in biodiesel production modeling. In this case, increasing levels of noise in the training data test the ability of the two different approaches to extract the true underlying behavior, and evaluate the robustness of KAN-ODEs to noisy data in light of the recent work suggesting the weakness of KANs in the presence of noise [30]. Second, we demonstrate ChemKANs as an efficient and time-saving surrogate model in an even larger system by learning zero-dimensional hydrogen combustion behavior using homogeneous reactor data that is subject to stiff dynamics, in a study designed to facilitate direct comparison against the MLP-based ChemNODE structure [17] that this second ChemKAN application is modeled after. In contrast to ChemNODE, where a reduced subset of the thermochemical state (excluding H, HO2, and H2O2) was learned using separately trained, non-interacting networks, we learn all species and temperature profiles here with a single compact ChemKAN network while retaining similar computational acceleration and performance. A successful demonstration of these tasks represents both an advancement in efficient and expressive machine learning modeling techniques for combustion, and an extension of KAN-ODEs as ChemKANs in systems significantly larger and more challenging than any previous KAN-ODE targets.
2 Methods
@@
This section describes the mathematical details in the newly developed chemistry Kolmogorov-Arnold network ordinary differential equations (ChemKANs). We begin with a review of a chemical kinetic model in Sec. 2.1, followed by existing MLP-based kinetic modeling techniques in Sec. 2.2. Then, we discuss the implementation of the ChemKAN framework for multi-purpose chemical model inference and computational acceleration in Sec. 2.3. That subsection includes a physics-enforced ChemKAN architecture, optimal learning strategies, and physics-informed loss functions. Finally, we provide the kinetic models of biodiesel pyrolysis and hydrogen-air combustion for demonstration of ChemKANs in Sec. 2.4.
2.1 Chemical kinetics model
@@
Given the state variables where is the temperature, is the mass fraction, and is the number of species, the net production/consumption rate for each species can be expressed as
@@
where is the time, is the density, is the molecular weight, and is the molar production or consumption rate [31]. While the specific functional form may differ across systems, is typically a strong function of temperature (for example, in the Arrhenius form ). In some cases, Eq. 1 sufficiently describes the chemical process where heat release or consumption is negligible. However, in other cases such as combustion and pyrolysis processes, chemical reactions entail exothermic and endothermic behaviors. In these cases, the temperature of a system can be tracked with energy conservation as in Eq. 2.
@@
where is the mixture-averaged specific heat and is the enthalpy of species . Here, the strongly nonlinear coupling of Y and T in Eqs. 1 and 2 often leads to modeling challenges such as numerical stiffness. We can express Eqs. 1-2 as a generic system of equations such that
@@ 2.2 MLP-based models
Recent machine learning approaches for model inference and computational acceleration often rely on neural networks constructed from Multi-layer perceptrons (MLPs). Among them, we introduce two mainstream models based on deep operator networks and neural ordinary differential equations. We will use these two MLP-based models to highlight the performance of our ChemKANs (to be introduced in Sec. 2.3).
2.2.1 Deep Operator Network (DeepONet)
@@
As one of the more popular architectures for combustion applications [21, 14, 22, 18], DeepONets [20] learn the chemical kinetic system (outlined in Sec. 2.1) through a physics-inspired separation between the system’s parameterization and the solution coordinate (Fig. 1(A)). Specifically, the solution is learned using two neural networks (branch net for thermochemical states and trunk net for time ), as per
@@
More specifically, given an initial condition , the DeepONet reports the solution at a given time as the element-wise product () of the branch network MLP evaluated on the initial condition and the trunk network MLP evaluated on the current time, with an optional final MLP layer for additional nonlinear encoding. are the learnable parameters for the respective neural networks. The DeepONet here, as well as in other augmented structures from higher-complexity implementations such as in [22] or [18]), directly learns the solution state at future times. By eliminating the integration step, these methods have shown substantial acceleration of computational times.
2.2.2 ChemNODEs
@@
In contrast to the DeepONet approach, ChemNODE [17] aims to learn the source terms in Eq. 3, rather than the direct thermochemical states over time . However, training challenges arose [17] when testing a single MLP network with inputs and outputs ( for and 1 for ). This led to development of a segregated model for each thermochemical state , as per
@@
Therefore, ChemNODE constructs MLP networks, each of which reads the entire current thermochemical state vector . The network has its own unique set of parameters , and is responsible for computing the current temporal gradient of that scalar . Notably, ChemNODE is still trained via MSE loss computed against the integrated solution profile (rather than the gradients that the MLPs directly output), thanks to the differentiatble Neural ODE framework [19]. This latent-dynamics modeling for allows ChemNODEs to compute rate terms without restriction to a specific instance. This contrasts with DeepONets which require an initial condition fed into the branch network. Therefore, ChemNODEs can be effortlessly generalized as the interpretable source terms even for large-scale simulations.
@@
We remark again that the split networks in this framework appear inefficient. In fact, this multi-network approach requires an -step training process, where each network is trained with the remaining thermochemical scalars held frozen. While this separated training strategy facilitates model convergence (and in fact was found necessary in Ref. [17] to converge the MLP gradient getters), it increases training cost exponentially with the number of species, especially when considering that no knowledge is shared between networks, even for common reactants. See Fig. 1 for visualizations of these two existing methods, as well as a simplified depiction of the method we will propose in the next sections.
@@
2.3 Chemistry Kolmogorov-Arnold network ordinary differential equations (ChemKANs)
2.3.1 Vanilla KAN-ODEs

### v2

@@ Abstract
Efficient chemical kinetic model inference and application in combustion are challenging due to large ODE systems and widely separated time scales. Machine learning techniques have been proposed to streamline these models, though strong nonlinearity and numerical stiffness combined with noisy data sources make their application challenging. Here, we introduce ChemKANs, a novel neural network framework with applications both in model inference and simulation acceleration for combustion chemistry. ChemKAN’s novel structure augments the generic Kolmogorov Arnold Network Ordinary Differential Equations (KAN-ODEs) with knowledge of the information flow through the relevant kinetic and thermodynamic laws. This chemistry-specific structure combined with the expressivity and rapid neural scaling of the underlying KAN-ODE algorithm instills in ChemKANs a strong inductive bias, streamlined training, and higher accuracy predictions compared to standard benchmarks, while facilitating parameter sparsity through shared information across all inputs and outputs. In a model inference investigation, we benchmark the robustness of ChemKANs to sparse data containing up to 15 added noise, and superfluously large network parameterizations. We find that ChemKANs exhibit no overfitting or model degradation in any of these training cases, demonstrating significant resilience to common deep learning failure modes. Next, we find that a remarkably parameter-lean ChemKAN (344 parameters) can accurately represent hydrogen combustion chemistry, providing a 2 acceleration over the detailed chemistry in a solver that is generalizable to larger-scale turbulent flow simulations. These demonstrations indicate the potential for ChemKANs as robust, expressive, and efficient tools for model inference and simulation acceleration for combustion physics and chemical kinetics.
I Introduction
Chemical kinetic modeling underpins a wide range of scientific and engineering applications, from biological systems to energy conversion. In combustion, two long-standing challenges hinder effective kinetic modeling: the complexity and size of detailed reaction mechanisms [1], and the computational stiffness these systems introduce during numerical integration [2, 3]. Recent advances in machine learning offer new opportunities to address both challenges through data-driven model discovery and accelerated surrogate solvers [4].
@@ I Introduction
For kinetic model discovery, a variety of learning algorithms, model structures, and optimization approaches have emerged. The Chemical Reaction Neural Network approach [5], for example, is capable of inferring reaction networks and parameters from limited species trajectory or heat release data [6, 7, 8] by directly enforcing the Arrhenius and mass action laws in a neural network structure. The Sparse Identification of Nonlinear Dynamics (SINDy) approach is similarly capable of extracting models from experimental data by assuming various functional relationship building blocks and learning the precise forms needed to fit the data [9]. Further optimization and inverse modeling tools exist for other chemical kinetic inference problems [10, 11, 12], with a key piece of many physics-based model inference techniques being a certain (and often substantial) degree of prior knowledge of the governing equations, reaction pathways, and reactants. A critical need across all of these methods is robustness to noisy data and model uncertainty, conditions common in combustion kinetics [5, 13, 7].
@@
On the solver front, researchers have proposed methods for dimension reduction [6, 14] and computational acceleration [15, 16, 17, 18] to handle stiff, high-dimensional systems. For instance, Owoyele and Pal [17] recently proposed ChemNODE, a creative and high-performing tool that uses the neural ODE concept of Chen et al. [19] to replace a complete chemical kinetic model with a collection of neural networks, one for each tracked thermochemical quantity. By using these networks to directly link the current thermochemical state to the chemical source term with no other problem-specific treatment, computational acceleration was enabled in the as-studied homogeneous reactor while retaining the generalizability of the surrogate model to higher-dimensional reacting flows where such acceleration becomes significantly more meaningful. Other recent works leverage DeepONets [20] to directly learn stiff integrators using neural operators, either with problem-specific network structures [21, 14] or by mapping from the current state to the source term, similar to ChemNODE [22, 18], allowing for significant computational acceleration downstream.
These strengths all come with drawbacks, however. Owoyele and Pal [17], for example, found that while the neural ODE approach’s clever exploitation of the dynamical structure of chemical kinetic models can provide high accuracy, the nonlinearity inherent to such models creates a challenging inference problem for the underlying MLP layers. This led the authors to omit a handful of species (including the key H radical) and break the training up into multiple unique and likely redundant networks, rather than a single cohesive architecture. Similarly, DeepONet techniques are cheap to evaluate and capture steady state behavior well, but their accuracy can suffer in stiff regions of the data that the integrator (which with DeepONets must be inferred directly by the network, as they do not explicitly leverage existing ODE solvers) can learn to skip over without significant penalties. We thus find that despite these recent novel and productive efforts, the training of efficient surrogate models for combustion chemistry remains a challenging task with open questions due to stiff behavior in the solution profiles and numerical instability in the nonlinear training processes.
@@
Kolmogorov-Arnold Networks were proposed recently [23] as an alternative to multi-layer perceptrons (MLPs) for general neural network applications, where instead of learning weights and biases on fixed activation functions, the shapes and magnitudes of the activation functions themselves are learned via gridded basis function sums and products. This shift was proposed to increase neural convergence rates, accuracy, and generalization. Echoing the development of traditional MLPs, physics-informed KAN structures were proposed shortly after, where it was found that certain knowledge of physical laws embedded in the training process can help the KAN converge to a physically meaningful solution [24, 25, 26]. Similar developments have been studied where direct encoding of physical insights or specific geometries into novel KAN structures has shown significant promise, such as including physical symmetries for quantum architecture search [27] or irregular geometries for flow simulations [28]. The inference benefits of KANs were additionally demonstrated to extend to dynamical system modeling in the Kolmorogorov-Arnold Network Ordinary Differential Equations (KAN-ODEs) framework [29], where KANs replaced MLPs in the neural ODE algorithm [19], and have since been demonstrated in a variety of settings, including predator-prey dynamics, shock formation, complex equations, phase separation [29], personalized cancer treatment [30], and flashover prediction [31].
In short, KAN-ODEs leverage KAN networks as gradient getters, while maintaining standard ODE solvers to integrate the solution profiles as in a traditional numerical approach. KAN-ODEs were shown to retain all major KAN benefits while also accessing the dynamical system inference capabilities of the neural ODE framework [29], which would appear to lend them to efficient chemical kinetic system modeling. However, a few key questions remain. KAN-ODEs have so far only been tested in relatively small systems (up to two-dimensional state variables), making their applicability and performance in larger, practical combustion systems unknown. Additionally, KANs in general have been shown to suffer substantially when trained with noisy datasets [32]. While we theorize that their direct coupling to ODE integrators combined with their sparse parameterization and smooth activations should provide KAN-ODEs with strong robustness to noise regardless of previously reported issues in generic KANs, this has not yet been studied quantitatively.
@@
In this work, we aim to develop a chemistry KAN-ODE (ChemKAN) framework for chemical kinetic modeling by designing the Kolmogorov-Arnold network gradient getter to learn the spatially-invariant relationship between the current thermochemical state and the chemical source terms. In the operator splitting regime, such a surrogate can be directly coupled to existing CFD or machine learning-based flow solvers for multi-dimensional combustion simulations in arbitrary physical domains. Developed across two case studies of increasing complexity, the ChemKAN framework contains a physics-informed, two-stage training process that enforces the direct coupling between species production and heat release, and additionally contains a soft constraint for element conservation. We further stabilize the optimization problems by implementing forward sensitivity analysis. Stiff chemistry is fully resolved via an attached numerical ODE solver. We study two cases here to explicitly probe the key behaviors and gaps in the model inference and solver acceleration literature identified above, both of which are addressed with ChemKANs.
First, we demonstrate the capability of ChemKANs to extract realistic and multi-species models from synthesized experimental data in a comparison against DeepONets (DONs) in biodiesel production modeling. In this case, increasing levels of noise in the training data test the abilities of the two different approaches to extract the true underlying behavior, and evaluate the robustness of ChemKANs (and KAN-ODEs in general) to noisy data in light of the recent work suggesting the weakness of KANs in the presence of noise [32]. Second, we demonstrate ChemKANs as efficient and time-saving surrogate models in an even larger system by learning zero-dimensional hydrogen combustion behavior using homogeneous reactor data that are subject to stiff dynamics, in a study designed to facilitate direct comparison against the MLP-based ChemNODE structure [17] that this second ChemKAN application was inspired by. In contrast to ChemNODE, where a reduced subset of the thermochemical state (excluding H, HO2, and H2O2) was learned using separately trained, non-interacting networks, we learn all species and temperature profiles here with a single compact ChemKAN network while retaining similar computational acceleration and performance. Across these two distinct cases, we demonstrate the strong capability and robustness of ChemKANs as efficient and expressive tools for both modeling and inference in combustion chemistry.
II Methods
@@
This section describes the mathematical details in the newly developed chemistry Kolmogorov-Arnold network ordinary differential equations (ChemKANs). We begin with a review of a zero-dimensional chemical kinetic model without transport effects in Sec. II.1, followed by existing MLP-based kinetic modeling techniques in Sec. II.2. Then, we discuss the implementation of the ChemKAN framework for multi-purpose chemical model inference and computational acceleration in Sec. II.3. That subsection includes a physics-enforced ChemKAN architecture, optimal learning strategies, and physics-informed loss functions. Finally, in Sec. II.4 we provide the kinetic models of biodiesel pyrolysis and hydrogen-air combustion that are used to demonstrate the performance of ChemKANs.
II.1 Chemical kinetics model
@@
Given the state variables where is the temperature, is the mass fraction, and is the number of species, the net production/consumption rate for each species in a homogeneous reactor can be expressed as
where is the time, is the density, is the molecular weight of species i, and is the molar production or consumption rate [33]. In the commonly used operator splitting approach, this homogeneous reactor is also applicable to higher-dimensional, turbulent simulations. While the specific functional form may differ across systems, is typically a strong function of temperature (for example, in the Arrhenius form ) as well as the current species concentrations. Additionally, energy conservation can be modeled by tracking the system temperature as in Eq. 2.
@@
In some cases, Eq. 1 sufficiently describes a chemical process when heat release or consumption is negligible. In other cases, such as combustion and pyrolysis processes, chemical reactions entail exothermic and endothermic behaviors. In these cases, the temperature of a system can be tracked with energy conservation, as in
@@
where is the mixture-averaged specific heat and is the enthalpy of species . Here, the strongly nonlinear coupling of Y and T in Eqs. 1 and 2 often leads to modeling challenges such as numerical stiffness. While always present in the governing laws and essential to accurate modeling for processes such as combustion and pyrolysis, Eq. 2 can occasionally be neglected, such as when modeling the kinetic rates of an isothermal experiment. We can express Eqs. 1-2 as a generic system of equations such that
@@ II.2 MLP-based models
Recent machine learning approaches for model inference and computational acceleration often rely on neural networks constructed from Multi-layer perceptrons (MLPs). Among them, we introduce two mainstream models based on deep operator networks and neural ordinary differential equations. We will use these two MLP-based models to highlight the performance of our ChemKANs (to be introduced in Sec. II.3).
II.2.1 Deep Operator Network (DeepONet)
@@
As one of the more popular architectures for combustion applications [21, 14, 22, 18], DeepONets [20] learn the chemical kinetic system (outlined in Sec. II.1) through a physics-inspired separation between the system’s parameterization and the solution coordinate (Fig. 1(A)). Specifically, the solution is learned using two neural networks (branch net for thermochemical states and trunk net for time ), as per
@@
More specifically, given an initial condition , the DeepONet reports the solution at a given time as the element-wise product () of the branch network MLP evaluated on the initial condition and the trunk network MLP evaluated on the current time, with an optional final MLP layer for additional nonlinear encoding. are the learnable parameters for the respective neural networks. The DeepONet here, as well as in other augmented structures from higher-complexity implementations [22, 18], directly learns the solution state at future times. By eliminating the integration step, these methods have shown substantial acceleration of computational times.
II.2.2 ChemNODEs
@@
In contrast to the DeepONet approach, ChemNODE [17] aims to learn the source terms in Eq. 3, rather than the direct thermochemical states over time . However, training challenges were observed [17] when testing a single MLP network with inputs and outputs ( for and 1 for ). This led to the development of a segregated model for each thermochemical state , as per
@@
Therefore, ChemNODE constructs MLP networks, each of which reads the entire current thermochemical state vector . The network has its own unique set of parameters , and is responsible for computing the current temporal gradient of that scalar . Notably, ChemNODE is still trained via MSE loss computed against the integrated solution profile (rather than the gradients that the MLPs directly output), thanks to the differentiable Neural ODE framework [19]. This latent-dynamics modeling for allows ChemNODEs to compute rate terms without restriction to a specific instance. This contrasts with standard DeepONets, which require an initial condition fed into the branch network. Therefore, ChemNODEs can be effortlessly generalized as the interpretable source terms even for large-scale simulations.
@@
We remark again that the split networks in this framework appear inefficient. In fact, this multi-network approach requires an -step training process, where each network is trained with the remaining thermochemical scalars held frozen. While this separated training strategy facilitates model convergence (and in fact was found necessary [17] to converge the MLP gradient getters), it increases training cost significantly with the number of species, especially when considering that no knowledge is shared between networks, even for common reactants. See Fig. 1 for visualizations of these two existing methods, as well as a simplified depiction of the method we will propose in the next sections.
@@
II.3 Chemistry Kolmogorov-Arnold network ordinary differential equations (ChemKANs)
II.3.1 Vanilla KAN-ODEs